OpenStudy (anonymous):

Solve the following systems of equations using the graphing method. What type of system is it? Name the solution if there is one. x+y=5 -x+y=3 show graph

OpenStudy (anonymous):

Get them in slope-intercept then graph them.

OpenStudy (anonymous):

???^^

OpenStudy (anonymous):

i still dont get it!

OpenStudy (anonymous):

Do you want me to show you?

OpenStudy (anonymous):

Yes, Please

OpenStudy (anonymous):

Do you know how to change an equation from standard form to slope-intercept?

OpenStudy (anonymous):

No, Iam only in 7th grade

OpenStudy (anonymous):

Well we are looking to get y by itself and to do that we have to move the x. So lets begin with x+y=5 take x and subtract it from both sides.

OpenStudy (anonymous):

Your equation is now y=-x+5 and is in slope-intercept.

OpenStudy (anonymous):

Can you just give me the answer??? cus iam tired!

OpenStudy (anonymous):

I suppose

OpenStudy (anonymous):

Hold on

OpenStudy (anonymous):

kk, Thx!

OpenStudy (anonymous):

what is the graph suppose to look like?

OpenStudy (anonymous):

Hold on I am making it for you now

OpenStudy (anonymous):

okie pokie dokie.....

OpenStudy (anonymous):

Should look like this It's independent or consistent

OpenStudy (anonymous):

ok thx,

OpenStudy (anonymous):

so you have y=-x+5 and then you still have to solve the other equation?

OpenStudy (anonymous):

-x+y=3

OpenStudy (anonymous):

we have to get y on one side

OpenStudy (anonymous):

correct?

OpenStudy (anonymous):

Correct, try the second one on your own

OpenStudy (anonymous):

it seems simple just add x to both sides then you have y=x+3

OpenStudy (anonymous):

correct?

OpenStudy (anonymous):

That's right, yeah

OpenStudy (anonymous):

awesome! thanks for the help on the graph you are really helpful

OpenStudy (anonymous):

Anytime :)